### v1

@@ Conclusion
By combining MD simulations with CNT analysis, we explore the kinetic competition between bulk, surface and defect cavitation of water under negative pressure. For defect-free surfaces, the prevalent cavitation pathway shifts abruptly from bulk to surface when the surface contact angle increases above a threshold in the range $\theta^{*}\approx 50^{\circ}$ to $60^{\circ}$, depending slightly on system size and waiting time. The typical cavitation pressure decreases in magnitude from about $p_{\mathrm{cav}}\approx-100$ MPa for bulk to about $p_{\mathrm{cav}}\approx-30$ MPa for hydrophobic surface cavitation. Even the presence of a single nanoscopic hydrophobic surface defect dominates the cavitation kinetics and substantially raises the cavitation pressure values, depending on the defect contact angle and defect size. Since even with the most advanced fabrication techniques surface defects are unavoidable, this explains why experimental cavitation pressures for typical surface materials are far above the predicted value for bulk cavitation and why low cavitation pressures around $p_{\mathrm{cav}}\approx-100$ MPa are only achieved in micro-sized defect-free quartz cavities (15, 16, 17).
In every-day life, water contains a variety of organic and inorganic substances. The bulk–surface cavitation crossover contact angle $\theta^{*}$ is close to the ‘Berg limit’ (47, 48, 49): Surfaces with larger contact angles are prone to macromolecule adhesion, which tends to make them more hydrophilic (50) and thus can prevent surface cavitation. Such a mechanism might be important for preventing embolies in the sap of tall trees. We have treated negative-pressure-induced cavitation in water at constant temperature. However, our methods can also be used for other liquids and to study cavitation in superheated liquids and nucleation in supersaturated solutions, where we expect surface and defect nucleation to play equally important roles.
[…]
Simulation model
The simulation setup follows our previous approach (44, 42). We model a planar surface composed of two self-assembled monolayers, each consisting of ten-carbon-atom alkyl chains terminated by polar hydroxyl ($\mathrm{OH}$) head groups. The molecules are arranged on a hexagonal lattice with a density of $4.3\,\mathrm{nm^{-2}}$. To stabilize the structure, we apply harmonic restraints to the second carbon atom from the head group with strengths of $k_{x}=k_{y}=500\,\mathrm{(kJ/mol)\,nm^{-2}}$ in the lateral directions, and to the tenth (terminal) carbon atom with $k_{x}=k_{y}=10\,\mathrm{(kJ/mol)\,nm^{-2}}$ laterally and $k_{z}=100\,\mathrm{(kJ/mol)\,nm^{-2}}$ in the normal direction. The surface polarity and thereby the surface contact angle is controlled by scaling the partial charges of the head groups with a factor $\alpha$, ranging from 0 to 1. For $\alpha=0$, the surface is non-polar with head groups resembling methylated termini, resulting in a contact angle of $\theta=135^{\circ}$. For $\alpha=1$, the surface is fully polar, mimicking hydroxyl head groups. The relationship between $\alpha$ and $\cos\theta$ is discussed in SI Appendix, section S4. The periodic simulation box, with initial dimensions $9\,\mathrm{nm}\times 10.4\,\mathrm{nm}\times 8\,\mathrm{nm}$, contains 16,353 water molecules and 800 hydroxylated alkanes.
We use united-atom parameters from the GROMOS force field for the alkanes (51, 42) and the SPC/E water model (52). Classical non-polarizable molecular dynamics simulations are performed with the GROMACS package (53) using a time step of $2\,\mathrm{fs}$. A plain cutoff of $0.9\,\mathrm{nm}$ is used for short-range Lennard-Jones interactions. Electrostatics is treated with the Smooth Particle–Mesh Ewald method, using a $0.9\,\mathrm{nm}$ real-space cutoff. A temperature of 300 K is controlled using the velocity rescale thermostat with a stochastic factor (54) and a time constant of $0.5\leavevmode\nobreak\ \mathrm{ps}$. Pressure is controlled by the Berendsen barostat with a time constant of $1\,\mathrm{ps}$, suitable even for negative pressures because of its efficiency and stability in box scaling under large pressure differences. We previously verified (35) that equivalent results are obtained with the Parrinello–Rahman barostat, which more accurately reproduces volume fluctuations. Reference coordinates for surface-atom restraints are scaled with the coupling matrix of the pressure coupling.
To simulate pressure ramps, we conduct a sequence of simulations at constant pressures, starting from an initial pressure of $-30\,\mathrm{MPa}$. After each simulation, the final coordinates are used to start the next simulation, with the pressure reduced by $0.1\,\mathrm{MPa}$. The duration of each constant-pressure simulation depends on the pressure rate $\dot{p}$.
Kinetic theory for pressure ramps
To obtain the cavitation attempt frequencies, we use the pressure-ramp simulation protocol (35), where the negative pressure decreases linearly over time as $p(t)=\dot{p}t$ at a constant rate $\dot{p}<0$. We identify cavitation when the simulation box length $L_{z}$ increases by more than 50 % of its initial value. The corresponding pressure-ramp cavitation pressure $p^{*}_{\mathrm{cav}}$ follows from the solution of the linear cavitation rate-equation as
$p^{*}_{\mathrm{cav}}=\dot{p}\int_{0}^{\infty}e^{-k_{0}I(t)}\mathrm{d}t\,,$ (16)
where
[…]
Here, $G^{*}(t)$ is the time-dependent free energy barrier in the pressure ramp protocol which, according to CNT, decay inversely with time as $G^{*}(t)\sim t^{-2}$. The attempt frequencies for bulk and surface cavitation, $\kappa_{\mathrm{3D}}$ and $\kappa_{\mathrm{2D}}$, are obtained from fits of Eq. 16 to the simulation data; detail are given in SI Appendix, section S5.
[…]
@@
Data, Materials, and Software Availability
All simulation input files as well as Python functions to fit $\kappa_{\mathrm{2D}}$ and perform further analysis from the simulations are available online.
[…]

### v2

@@ IV Conclusion
By combining MD simulations with the CNT framework, we explore the kinetic competition among the three fundamental pathways of cavitation in water under negative pressure: bulk cavitation, surface cavitation and defect-induced cavitation. For defect-free surfaces, the prevalent cavitation pathway shifts abruptly from bulk to surfaces when the surface contact angle increases above a threshold of $\theta^{*}\approx 50^{\circ}$ to $60^{\circ}$, depending slightly on system size and waiting time. The typical cavitation pressure decreases in magnitude from about $p_{\mathrm{cav}}\approx-100$ MPa for bulk to about $p_{\mathrm{cav}}\approx-30$ MPa for hydrophobic surface cavitation. Even the presence of a single nanoscopic hydrophobic surface defect dominates the cavitation kinetics and substantially raises the cavitation pressure, depending on the defect contact angle and defect size. Since even with the most advanced fabrication techniques surface defects are unavoidable and experimental cavitation pressures for typical surface materials are far above the predicted value for bulk cavitation and low cavitation pressures around $p_{\mathrm{cav}}\approx-100$ MPa are only achieved in micro-sized defect-free quartz cavities Zheng et al. (1991); Alvarenga, Grimsditch, and Bodnar (1993); Azouzi et al. (2013b).
While hydrophobic aggregates reduce the water stability against cavitation, individual dispersed molecules generally have a minimal effect Zhou et al. (2019); Kanduč et al. (2020). Viscosifying agents, such as polymers, slow down cavitation kinetics, by influencing the prefactor $\kappa$ in Eqs. (6) and (11) without altering the exponential term. However, surfaces with contact angles exceeding the “Berg limit” of approximately 65°—coincidentally near the crossover contact angle—are prone to adhesion of amphiphilic molecules Vogler (1998); Rosenhahn et al. (2010); Šako et al. (2023); Kanduč, Schneck, and Netz (2024). This adhesion can render such surfaces more hydrophilic and coat surface defects Šako et al. (2023), thereby suppressing surface cavitation. This mechanism might explain the prevention of embolisms through lipid adsorption in tree sap Schenk, Steppe, and Jansen (2015) or the increased cavitation stability observed when amphiphilic polymers are added to water Brujan et al. (2008); Gruzdkov and Petrov (2008). Exploring the effect of molecular adsorption on defects and cavitation is an intriguing question, which we aim to pursue in upcoming studies.
In this study, we focused on pressure-induced cavitation in water at constant temperature. However, the methods developed here can be extended to other liquids and applied to cavitation in superheated liquids or nucleation in supersaturated solutions. In such systems, surface and defect nucleation are expected to play similarly critical roles.
V Supplementary Material
The supplementary material contains additional information on fitting the kinetic prefactor, the cavitation pressure for various waiting times $\boldsymbol{\tau}$, the relation between surface polarity $\alpha$ and contact angle of the SAM surface.
Acknowledgements.
@@
Data Availability Statement
All simulation input files as well as Python functions to fit $\kappa_{\mathrm{2D}}$ and perform further analysis from the simulations are available on Zenodo.